OpenStudy (anonymous):

find the slope intercept equation slope 0 and y-intercept (0,4) write in y=mx+b form

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

slope is zero, means the line looks like \(y=b\) in this case \(b=4\)

OpenStudy (anonymous):

so how would i solve it though?

OpenStudy (anonymous):

you don't "solve" it, you write it \[y=4\]

OpenStudy (nikato):

That's one way to think of it. Or if you can't see that or understand it, you can always plug in numbers. y=mx+b We know m=0. x=0. y=4 4=0(0)+b So b=4 Put that back in the equation. Plug b and m in only. y= 0(x)+4. Remember any number times 0 always =0 So y=4